OpenStudy (anonymous):

If y varies inversely as x, and y = 10 when x = 7, find y for the x-value of 10. 3 2 6 7

OpenStudy (mathhelp346):

when something is inversely related : y=k/x DONT FORGET THE CONSTANT k :D now lets solve for k (since its a constant, it ALWAYS STAYS THE SAME :D) 10=k/7, thus k=70 now lets make the second and final equation, y=70/10 thus, y=7

OpenStudy (anonymous):

thank youu!:)

OpenStudy (mathhelp346):

np :)